OpenStudy (anonymous):

which value of x is the solution of 2x/5 + 1/3=7x-2/15

OpenStudy (anonymous):

You have to solve for X this means that all x have to be on one sides. For example you want 7x to be on the left. SO you will have to subtract both sides by 7x 2x/5 + 1/3=7x-2/15 2x/5 -7x + 1/3=7x-7x-2/15 2x/5 -7x + 1/3=-2/15 Now you do that rest. Get rid of that 1/3 from the left side.

OpenStudy (anonymous):